# A geometric dual of F-maximization in massive type IIA

### v1

@@ 2 Review of equivariant localization
where $\epsilon^{i}$ are the weights of $\xi$ on $N_{\Sigma}$, i.e. in local coordinates $\xi=\sum_{i=1}^{k}\epsilon_{i}\partial_{\varphi_{i}}$ with $\partial_{\varphi_{i}}$ rotating $L_{i}$, and $c_{1}(L_{i})$ are the first Chern classes. In practice the denominator is to be expanded in a power series in the Chern class, as illustrated below, which truncates at order dim$(\Sigma)$ at most. In summary, the game to be played is to build equivariantly closed polyforms whose top form is a physical quantity that we want to measure. The physical observable, obtained by integrating the top form over some (sub-)manifold of our internal space, is then evaluated using the BV–AB formula. This gives the observable as a sum of contributions of the lower forms evaluated on fixed submanifolds.
@@
As we will see shortly, in our setup the internal manifold can have boundaries and this requires an extension of the BV–AB formula (2.3). We make the restriction that the boundary of the manifold does not have any fixed points and let $f:\partial M\hookrightarrow M$ be the embedding of the boundary into $M$. Then to the BV–AB formula in (2.3) we add in an additional boundary contribution giving:
@@
$\frac{1}{d_{\xi}\xi}=-\frac{1}{|\xi|^{2}(1-|\xi|^{-2}d\xi)}=-\frac{1}{|\xi|^{2}}\bigg{[}1+\sum_{a=1}^{\infty}\left(\frac{\mathrm{d}\xi}{|\xi|^{2}}\right)^{a}\bigg{]}\,.$ (2.6)
It is obvious from the formula that the boundary must not have fixed points of the U$(1)$ given the factors of $|\xi|^{2}$ appearing in the denominator.
It turns out that the boundary contributions from the observables we will consider vanish because $f^{*}\Phi=0$ for our integrals. This is not a generic feature and is special to the various setups that we consider. One expects that by looking at different admissible boundary conditions or more refined observables, boundary contributions will be important. We hope to study this in the future.
@@ 3.1 Setup
The above fluxes are the (magnetic) ones which appear in the equations of motion. They, however, are not the ones one would use to quantize the fluxes. These are the Page fluxes: $f=F\wedge\mathrm{e}^{-B}$ in polyform notation. Here $F$ is the polyform of magnetic fluxes, $B$ is either of the representatives of $B$ given in (3.9) and one extracts out the the part of form degree $p$. Recall that Page fluxes are closed, hence define conserved charges, but are not gauge invariant. The quantization condition reads
@@ 3.2 Degenerations and global regularity
We will be interested in capping off the spacetime with an O8-plane singularity. In general the metric and dilaton behaviour of such a singularity is
@@
where we have kept only the leading order pieces and the plane is located at $r=0$.
@@ 3.3.1 The $l\neq 0$ case
After a slightly tedious, but straightforward computation, one can compute call the various polyforms outlined above. We suppress the details of the computations, presenting just the final polyforms. The polyform for the free energy is
@@ 3.4 Localization
$Y^{\prime}=-2\Phi_{2}^{F_{4}}\,.$ (3.30)
The BV–AB formula then gives
@@ 4.1 D4-branes wrapped on a Riemann surface
and in terms of these variables the free energy reads:
@@
$\mathcal{F}=\frac{2\pi\kappa(1-g)N^{5/2}}{5n_{0}^{1/2}}(\sqrt{\kappa^{2}+8z^{2}}-3\kappa)\sqrt{1-\frac{\kappa(\sqrt{\kappa^{2}+8z^{2}}+\kappa)}{4z^{2}}}\,.$ (4.13)
@@
$\Delta(\Sigma_{g}^{p})=-\frac{1}{(2\pi)^{2}\ell_{s}^{3}}\int_{\Sigma_{g}^{p}}\frac{1}{F_{0}y}j=-2(b_{1}p_{2}+b_{2}p_{1})N=2(1-g)(z\varepsilon-\kappa)N\,.$
@@
$\begin{split}\Delta(HS^{2,i})&=\left(1\mp\frac{\sqrt{\kappa^{2}+8z^{2}}+\kappa}{4z}\right)N\,,\\ \Delta(\Sigma_{g}^{p})&=\frac{(1-g)}{2}\left(\sqrt{\kappa^{2}+8z^{2}}-3\kappa\right)N\,,\end{split}$ (4.15)
with the $\mp$ signs corresponding to $i=1,2$ respectively. The on-shell expressions match Bah:2018lyv . Note that requiring these to be positive fixes $|z|>1$ for $\kappa=1$ and $|z|>0$ for $\kappa=-1$. This is indeed the correct field theory constraints and it is satisfying that it pops out from gravity.
@@
$\begin{split}\Delta(HS^{2,i})&=\left(1\mp\frac{\kappa-\sqrt{\kappa^{2}+8z^{2}}}{4z}\right)N\,,\\ \Delta(\Sigma_{g}^{p})&=\left(\frac{\kappa(g-1)(3\kappa+\sqrt{\kappa^{2}+8z^{2}})}{2}\right)N\,.\end{split}$ (4.16)
@@ 4.2 D4-branes wrapped on a spindle
An obvious obstruction to finding explicit metrics is to require these conformal dimensions to be positive, however it is clear that since the $y_{\pm}$ always appear squared this will not fix the sign ambiguity noted in footnote 12. We leave understanding these points to future work.
@@ 4.4 $S^{2}$ bundle over $B_{4}$
$c_{\alpha}^{S}-c_{\alpha}^{N}=2l\ell_{s}\pi k_{2}n_{\alpha}\,,\quad\Leftrightarrow\quad k_{\alpha}^{S}-k_{\alpha}^{N}=k_{2}n_{\alpha}\,.$ (4.52)
@@
$\mathcal{F}=\frac{4\pi}{3}\sqrt{\frac{2k_{2}^{3}L^{3}(L-2)^{3}}{\langle k^{N},k^{N}\rangle L^{2}+\langle k^{S},k^{S}\rangle(L-2)^{2}+\langle k^{N},k^{S}\rangle L(L-2))}}N^{3/2}\,.$ (4.57)
@@ 4.5.1 Round five-sphere
$0=\int_{S_{i}^{2}}\Phi^{Y}=-\frac{\pi}{3b_{i}}\left[\frac{1}{y_{N}}(F_{0}^{2}y_{N}^{4}-3l^{2})-\frac{1}{y_{S}}(F_{0}^{2}y_{N}^{S}-3l^{2})\right]\,,$ (4.66)
@@ 4.5.2 Quasi-regular Sasaki–Einstein
We now want to understand what $\mathcal{E}$ is computing. Let us take $\mathcal{N}_{N}=-\mathcal{N}_{S}$ for simplicity. Then for a Calabi–Yau cone we have that [^19] One can make this more precise by using the results of Martelli:2006yb ; goertsches2015localization ; Boyer_2018 . One can understand the factors of $3$ due to the same normalization of the Reeb vector field in the previous $S^{5}$ example, 3 of the 2’s due to the formula for the volume and one factor of the 2’s since $\beta$ double counts the volume when $\mathcal{N}_{N}=\mathcal{N}_{S}$.
@@ 5 Conclusions and future directions
So far all the supergravity equivariantly localized solutions have used an even-dimensional localizing space. Notice that in section 4.4 we have actually localized over an odd-dimensional space. The construction uses an additional circle to first reduce on and produce an even-dimensional space over which one can now localize. There are various tools, beyond this reduction, that one may use to localize over odd-dimensional spaces Boyer:2017jwa ; goertsches2015localization . It would be interesting to extend these techniques to these odd-dimensional examples. Moreover one could further extend to including higher derivative corrections and therefore go beyond the large $N$ limit from supergravity. We hope to return to some of these future directions soon.
@@ Appendix A Some cohomology and homology
We may define the Thom class for the bundle. Let $\rho(r)$ be a radial function which is $1$ at the centre of the disc and 0 on the boundary. This then makes $\mathrm{d}\rho=\rho^{\prime}(r)\mathrm{d}r$ a bump form on $\mathbb{R}$ with total integral 1. Further, define the global angular form on the disc to be $\eta$. Then the Thom class may be written as212121 We thank Topology Boy for clarifying this and other points in this appendix with us.

### v2

@@ 2 Review of equivariant localization
where $\epsilon_{i}$ are the weights of $\xi$ on $N_{\Sigma}$, i.e. in local coordinates $\xi=\sum_{i=1}^{k}\epsilon_{i}\partial_{\varphi_{i}}$ with $\partial_{\varphi_{i}}$ rotating $L_{i}$, and $c_{1}(L_{i})$ are the first Chern classes. In practice the denominator is to be expanded in a power series in the Chern class, as illustrated below, which truncates at order dim$(\Sigma)$ at most. In summary, the game to be played is to build equivariantly closed polyforms whose top form is a physical quantity that we want to measure. The physical observable, obtained by integrating the top form over some (sub-)manifold of our internal space, is then evaluated using the BV–AB formula. This gives the observable as a sum of contributions of the lower forms evaluated on fixed submanifolds.
@@
2.1 Boundary contributions
As we will see shortly, in our setup the internal manifold can have boundaries and this requires an extension of the BV–AB formula (2.3). We make the restriction that the boundary of the manifold does not have any fixed points and let $f:\partial M\hookrightarrow M$ be the embedding of the boundary into $M$. Then to the BV–AB formula in (2.3) we add in an additional boundary contribution Szabo:1996md :
@@ 2.1 Boundary contributions
$\frac{1}{\mathrm{d}_{\xi}\xi}=-\frac{1}{|\xi|^{2}(1-|\xi|^{-2}\mathrm{d}\xi)}=-\frac{1}{|\xi|^{2}}\bigg{[}1+\sum_{a=1}^{\infty}\left(\frac{\mathrm{d}\xi}{|\xi|^{2}}\right)^{a}\bigg{]}\,.$ (2.6)
It is obvious from the formula that the boundary must not have fixed points of the U$(1)$ given the factors of $|\xi|^{2}$ appearing in the denominator. This was a choice that we made and there exists a more general formula when this is not true, Szabo:1996md .
To better explain the contributions from the boundary let us study an example. Consider a four-sphere and let us put in an arbitrary boundary which preserves a choice of U$(1)$ action that we will specify shortly. Let us take the following metric on the four-sphere [^4] We use the same conventions as in BenettiGenolini:2023ndb in order to use their polyforms.
$\mathrm{d}s^{2}=\mathrm{d}\alpha^{2}+\sin^{2}\alpha\big{(}\mathrm{d}\theta^{2}+\sin^{2}\theta\mathrm{d}\phi_{1}^{2}+\cos^{2}\theta\mathrm{d}\phi_{2}^{2}\big{)}\,.$ (2.7)
For the round four-sphere the ranges of the coordinates are: $\alpha\in[0,\pi]$, $\theta\in[0,\tfrac{\pi}{2}]$, and $\phi_{1},\phi_{2}\in[0,2\pi]$. We will localize using the Killing vector
$\xi=b_{1}\partial_{\phi_{1}}+b_{2}\partial_{\phi_{2}}\,,$ (2.8)
with both $b$’s non-zero. The norm of this Killing vector field is
$|\xi|^{2}=\sin^{2}\alpha(b_{1}^{2}\sin^{2}\theta+b_{2}^{2}\cos^{2}\theta)\,,$ (2.9)
which clearly vanishes at the poles of the $S^{4}$ at $\alpha=0,\pi$. Rather than taking the round four-sphere we will restrict $\alpha$ to $\alpha\in[0,\alpha_{0}]$ with $0<\alpha_{0}<\pi$ and call the space $X_{\alpha_{0}}$. For $\alpha_{0}=\tfrac{\pi}{2}$ this is the usual four-dimensional hemi-sphere. It is simple to explicitly compute the volume for this metric and we find
$\mathrm{Vol}(X_{\alpha_{0}})=\frac{4\pi^{2}}{3}+\frac{2\pi^{2}\cos\alpha_{0}(\cos\alpha_{0}^{2}-3)}{3}\,.$ (2.10)
We now want to reproduce this using our localization formulae. We have a fixed point at the pole and a boundary term at $\alpha=\alpha_{0}$. Using the results in BenettiGenolini:2023ndb the equivariantly closed polyform for the (weighted) volume form is given by $\Phi=\Phi_{4}+\Phi_{2}+\Phi_{0}$ with:
$\begin{split}\Phi_{4}&=\frac{1}{b_{1}b_{2}}\mathrm{vol}=\frac{1}{b_{1}b_{2}}\sin^{3}\alpha\sin\theta\cos\theta\mathrm{d}\alpha\wedge\mathrm{d}\theta\wedge\mathrm{d}\phi_{1}\wedge\mathrm{d}\phi_{2}\,,\\ \Phi_{2}&=-\frac{1}{2}\sin^{3}\alpha\left(\frac{1}{b_{1}}\sin^{2}\theta\mathrm{d}\phi_{1}+\frac{1}{b_{2}}\cos^{2}\theta\mathrm{d}\phi_{2}\right)\wedge\mathrm{d}\alpha\,,\\ \Phi_{0}&=\frac{1}{6}(3\cos\alpha-\cos^{3}\alpha)\,.\end{split}$ (2.11)
The weighted volume is then
$\begin{split}\int_{X_{\alpha_{0}}}\frac{1}{b_{1}b_{2}}\mathrm{vol}(X_{\alpha_{0}})&=\frac{(2\pi)^{2}}{b_{1}b_{2}}\Phi_{0}\Big{|}_{\alpha=0}+\int_{\partial X_{\alpha_{0}}}f^{*}\frac{1}{|\xi|^{2}}\left[\xi\wedge\Phi_{2}+\frac{\Phi_{0}}{|\xi|^{2}}\xi\wedge\mathrm{d}\xi\right]\\ &=\frac{4\pi^{2}}{3b_{1}b_{2}}+\int_{\partial X_{\alpha_{0}}}\frac{b_{1}b_{2}\cos\alpha_{0}(3-\cos^{2}\alpha_{0})\cos\theta\sin\theta}{3(b_{1}^{2}\sin^{2}\theta+b_{2}^{2}\cos^{2}\theta)^{2}}\mathrm{d}\theta\wedge\mathrm{d}\phi_{1}\wedge\mathrm{d}\phi_{2}\end{split}$ (2.12)
Note that the contribution only arises from the second term involving $\Phi_{0}$ since $f^{*}\Phi_{2}=0$. We now need to perform the integral over $\partial X_{\alpha_{0}}$. Note that this is just an integral over a round three sphere and one could in principle use equivariant localization again to perform this integral. We will refrain from doing this but it is not difficult to see that the final result is:
$\int_{X_{\alpha_{0}}}\frac{1}{b_{1}b_{2}}\mathrm{vol}(X_{\alpha_{0}})=\frac{2\pi^{2}(2-3\cos\alpha_{0}+\cos^{3}\alpha_{0})}{3b_{1}b_{2}}\,,$ (2.13)
and therefore we find that we recover the correct result in (2.10). It is important to note that for generic $\alpha_{0}\neq\tfrac{\pi}{2}$ there is a contribution from the boundary. One expects that it is possible to use equivariant localization to also evaluate the integrals for the boundary contributions and this would be an interesting problem to study in the future.
Having given the generic formula with boundaries it turns out that the boundary contributions from the observables we will consider vanish because $f^{*}\Phi=0$ for our integrals. This is not a generic feature, indeed the example above had a boundary contribution, except if we looked at the hemi-sphere. Rather, it is special to the various setups that we consider and the types of boundary there. One expects that by looking at different admissible boundary conditions or more refined observables, boundary contributions will no longer vanish. We hope to study this in the future.
@@ 3.1 Setup
The above fluxes are the (magnetic) ones which appear in the equations of motion. They, however, are not the ones one would use to quantize the fluxes. These are the Page fluxes: $f=F\wedge\mathrm{e}^{-B}$ in polyform notation. Here $F$ is the polyform of magnetic fluxes, $B$ is either of the representatives of $B$ given in (3.9) and one extracts out the part of form degree $p$. Recall that Page fluxes are closed, hence define conserved charges, but are not gauge invariant. The quantization condition reads
@@ 3.2 Degenerations and global regularity
We will be interested in capping off the spacetime with an O8-plane singularity with divergent dilaton. The metric and dilaton behaviour of such an O8 singularity is
@@
where we have kept only the leading order pieces and the plane is located at $r=0$. Note that there are other choices of O8-plane which would not have the above divergence structure, this choice is taken to match with the O8-plane in the Brandhuber–Oz solution Brandhuber:1999np that will be relevant later.
@@ 3.3.1 The $l\neq 0$ case
After a slightly tedious, but straightforward computation, one can compute all the various polyforms outlined above. We suppress the details of the computations, presenting just the final polyforms. The polyform for the free energy is
@@ 3.4 Localization
$Y^{\prime}=F_{0}\Big{[}\frac{y}{F_{0}^{2}\mathrm{e}^{4A}}j+\frac{y}{2}\mathrm{d}y\wedge D\psi\Big{]}=-2\Phi_{2}^{F_{4}}\,,$ (3.30)
with $\Phi_{2}^{F_{4}}$ the two-form part of the polyform for $F_{4}$ in equation (3.25). The BV–AB formula then gives
@@ 4.1 D4-branes wrapped on a Riemann surface
with $\kappa$ the extrinsic curvature $\kappa=1$ for $g=1$, $\kappa=0$ for $g=1$ and $\kappa=-1$ for $g>1$. In terms of these variables the free energy reads:
@@
$\mathcal{F}=\frac{2\pi(1-g)N^{5/2}}{5\kappa n_{0}^{1/2}}(\sqrt{\kappa^{2}+8z^{2}}-3\kappa)\sqrt{1-\frac{\kappa(\sqrt{\kappa^{2}+8z^{2}}+\kappa)}{4z^{2}}}\,.$ (4.13)
@@
$\Delta(\Sigma_{g}^{p})=-\frac{1}{(2\pi)^{2}\ell_{s}^{3}}\int_{\Sigma_{g}^{p}}\frac{1}{F_{0}y}j=-2(b_{1}p_{2}+b_{2}p_{1})N=\frac{2(1-g)}{\kappa}(z\varepsilon-\kappa)N\,.$
@@
$\begin{split}\Delta(HS^{2,i})&=\left(1\mp\frac{\sqrt{\kappa^{2}+8z^{2}}+\kappa}{4z}\right)N\,,\\ \Delta(\Sigma_{g}^{p})&=\frac{(1-g)}{2\kappa}\left(\sqrt{\kappa^{2}+8z^{2}}-3\kappa\right)N\,,\end{split}$ (4.15)
with the $\mp$ signs corresponding to $i=1,2$ respectively. The on-shell expressions match Bah:2018lyv . Note that requiring these to be positive fixes $|z|>1$ for $\kappa=1$ and $|z|>0$ for $\kappa=-1$. The torus case can be obtained by setting $\kappa=0$ and $g=1$ in both expressions with $\tfrac{1-g}{\kappa}\rightarrow 1$. This is indeed the correct field theory constraints and it is satisfying that it pops out from gravity.
@@
$\begin{split}\Delta(HS^{2,i})&=\left(1\mp\frac{\kappa-\sqrt{\kappa^{2}+8z^{2}}}{4z}\right)N\,,\\ \Delta(\Sigma_{g}^{p})&=\left(\frac{(g-1)(3\kappa+\sqrt{\kappa^{2}+8z^{2}})}{2\kappa}\right)N\,.\end{split}$ (4.16)
@@ 4.2 D4-branes wrapped on a spindle
An obvious obstruction to finding explicit metrics is to require these conformal dimensions to be positive, however it is clear that since the $y_{\pm}$ always appear squared this will not fix the sign ambiguity noted in footnote 13. We leave understanding these points to future work.
@@ 4.4 $S^{2}$ bundle over $B_{4}$
$c_{\alpha}^{S}-c_{\alpha}^{N}=2l\ell_{s}\pi k_{2}n_{\alpha}\quad\Leftrightarrow\quad k_{\alpha}^{S}-k_{\alpha}^{N}=k_{2}n_{\alpha}\,.$ (4.52)
@@
$\mathcal{F}=\frac{4\pi}{3}\sqrt{\frac{2k_{2}^{3}L^{3}(L-2)^{3}}{\langle k^{N},k^{N}\rangle L^{2}+\langle k^{S},k^{S}\rangle(L-2)^{2}+\langle k^{N},k^{S}\rangle L(L-2)}}N^{3/2}\,.$ (4.57)
@@ 4.5.1 Round five-sphere
$0=\int_{S_{i}^{2}}\Phi^{Y}=-\frac{\pi}{3\epsilon_{i}}\left[\frac{1}{y_{N}}(F_{0}^{2}y_{N}^{4}-3l^{2})-\frac{1}{y_{S}}(F_{0}^{2}y_{N}^{S}-3l^{2})\right]\,,$ (4.66)
@@ 4.5.2 Quasi-regular Sasaki–Einstein
We now want to understand what $\mathcal{E}$ is computing. Let us take $\mathcal{N}_{N}=-\mathcal{N}_{S}$ for simplicity. Then for a Calabi–Yau cone we have that [^20] One can make this more precise by using the results of Martelli:2006yb ; Goertsches:2015vga ; Boyer:2017jwa . One can understand the factors of $3$ due to the same normalization of the Reeb vector field in the previous $S^{5}$ example, 3 of the 2’s due to the formula for the volume and one factor of the 2’s since $\mathcal{E}$ double counts the volume when $\mathcal{N}_{N}=\mathcal{N}_{S}$.
@@ 5 Conclusions and future directions
So far all the supergravity equivariantly localized solutions have used an even-dimensional localizing space. Notice that in section 4.4 we have actually localized over an odd-dimensional space. The construction uses an additional circle to first reduce on and produce an even-dimensional space over which one can now localize. There are various tools, beyond this reduction, that one may use to localize over odd-dimensional spaces Boyer:2017jwa ; Goertsches:2015vga . It would be interesting to extend these techniques to these odd-dimensional examples. Moreover one could further extend to including higher derivative corrections and therefore go beyond the large $N$ limit from supergravity. We hope to return to some of these future directions soon.
@@ Appendix A Some cohomology and homology
We may define the Thom class for the bundle. Let $\rho(r)$ be a radial function which is $1$ at the centre of the disc and 0 on the boundary. This then makes $\mathrm{d}\rho=\rho^{\prime}(r)\mathrm{d}r$ a bump form on $\mathbb{R}$ with total integral 1. Further, define the global angular form on the disc to be $\eta$. Then the Thom class may be written as222222 We thank Topology Boy for clarifying this and other points in this appendix with us.